# Learning the Mechanism of Catastrophic Forgetting: A Perspective from Gradient Similarity

### v1

@@ 5.1 Establishment of CNL
To indicate the sign of gradient similarity for the $j$-th neuron, we define an indicator function $I(\cdot)$ as
@@
Applying the indicator function to all neurons yields
$\mathbb{I}[S(\mathcal{M},\mathcal{I})\geq 0]$ (18)
$=\big(I(s_{\theta_{1}}(\mathcal{M},\mathcal{I})\geq 0),\ldots,I(s_{\theta_{|\theta|}}(\mathcal{M},\mathcal{I})\geq 0)\big)^{\top}$
CNL freezes neurons with negative gradient similarity, leading to a parameter update rule defined as
@@
where $\odot$ denotes the Hadamard (element-wise) product. According to Eq. (2.2), the change in loss on Mastered Set $\mathcal{M}$ after parameter update is given by
$\Delta L^{\mathcal{M}}(\theta)$
$=-\eta\,\nabla_{\theta}L^{\mathcal{M}}(\theta)^{\top}\big(\mathbb{I}[S(\mathcal{M},\mathcal{I})\geq 0]\odot\nabla_{\theta}L^{\mathcal{I}}(\theta)\big)$
@@
In practical settings, it is almost impossible for all parameter gradients to be exactly zero. Therefore, we have
$\Delta L^{\mathcal{M}}(\theta)$ (21)
$=-\eta\sum_{j=1}^{|\theta|}\Big[I(s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0)\,s_{\theta_{j}}(\mathcal{M},\mathcal{I})\Big]\;<0$
This result indicates that, under the assumption of an infinitesimally small learning rate $\eta$, CNL guarantees a monotonic decrease of the loss on Mastered Set $\mathcal{M}$. In other words, CNL theoretically prevents catastrophic forgetting.
To experimentally evaluate the effectiveness of CNL, we performed knowledge injection using both CNL and conventional fine tuning (FT). Their forgetting was compared under the constraint that CNL learned at least as many new samples as FT. As shown in Table 3, FT forgets a number of mastered samples comparable to the amount of newly injected knowledge, whereas CNL achieve zero forgetting.
Although CNL exhibits excellent performance in preventing forgetting, we have to point out that CNL achieves zero forgetting only under two restrictive conditions: (i) the learning rate $\eta$ is sufficiently small, and (ii) Mastered Set $\mathcal{M}$ is explicitly known. The case where Mastered Set $\mathcal{M}$ is not explicitly available will be discussed in Section 5.4.
5.2 Extending CNL to More Complex Optimizers
Table 4: Comparison of knowledge injection performance across different optimizers. Qwen2.5 1.5B is selected for experiment. Whether using SGD, Momentum, Adam, or AdamW, CNL consistently achieves zero catastrophic forgetting while maintaining effective knowledge injection.
Optimizer Method MMLU MedQA ARC-C CSQA
@@ 5.2 Extending CNL to More Complex Optimizers
SGD FT 168 188 191 264 104 88 86 163
(24.8%) (21.4%) (27.1%) (46.5%) (32.1%) (9.3%) (31.6%) (19.8%)
CNL 168 0 193 0 163 0 147 0
(24.8%) (0.0%) (27.1%) (0.0%) (50.3%) (0.0%) (54.0%) (0.0%)
Momentum FT 341 450 257 398 261 169 205 233
(50.3%) (51.3%) (36.5%) (70.1%) (80.6%) (17.9%) (75.4%) (28.3%)
CNL 408 0 461 0 291 0 240 0
(60.2%) (0.0%) (65.5%) (0.0%) (89.8%) (0.0%) (88.2%) (0.0%)
Adam FT 647 203 704 181 320 143 266 168
(95.4%) (23.1%) (100.0%) (31.9%) (98.8%) (15.1%) (97.8%) (20.4%)
CNL 648 0 704 0 323 0 266 0
(95.6%) (0.0%) (100.0%) (0.0%) (99.7%) (0.0%) (97.8%) (0.0%)
AdamW FT 638 195 703 174 320 143 251 294
(94.1%) (22.2%) (99.9%) (30.6%) (98.8%) (15.1%) (92.3%) (35.7%)
CNL 642 0 703 0 323 0 266 0
(94.7%) (0.0%) (99.9%) (0.0%) (99.7%) (0.0%) (97.8%) (0.0%)
After establishing CNL with Stochastic Gradient Descent (SGD) optimizer Rumelhart et al. (1985), we further extend CNL to Momentum Rumelhart et al. (1985), Adam Kingma (2014), AdamW Loshchilov and Hutter (2019), and arbitrary optimizer.
In Appendix G, rigorous proofs are provided that the loss on Mastered Set decreases monotonically, as long as the parameter update of Momentum, Adam, AdamW, arbitrary optimizer follows Eq. (27), (38), (43), and (58).
Table 4 demonstrates the experimental results of CNL with difference optimizers. When FT is applied, all four optimizers suffer from severe catastrophic forgetting. In contrast, CNL effectively prevents catastrophic forgetting across all four optimizers. In particular, when using Adam optimizer, CNL is able to successfully inject nearly all (95.6%-100.0%) required knowledge while exhibiting zero forgetting.
Additionally, the dynamic characteristic of learning and forgetting is analyzed in Appendix. H. Although CNL freezes conflicting neurons and trains only collaborative neurons, its learning efficiency remains comparable to that of conventional FT. This further demonstrates the engineering potential of our CNL.
5.3 Extending CNL to More Complex Datasets
To further valid CNL on more complex datasets, MMLU, MedQA, ARC-C, and CSQA were merged into a unified dataset, referred to as MMAC. The performance of FT and CNL on MMAC is studied in Appendix I.
Results demonstrate that, while successfully injecting 95% knowledge into LLMs, FT leads to 91.7% forgetting of mastered knowledge, while CNL exhibits zero forgetting. These findings further demonstrate that CNL can effectively prevent catastrophic forgetting even under more complex and large-scale data conditions.
5.4 Extending CNL to Out-of-Set Generalization
We have verified that CNL achieves zero forgetting under in-set setting. However, it is hard to explicitly obtain all Mastered Set and Injection Set in practice. Therefore, the out-of-set generalization capability of CNL is analyzed.
First, Injection Set $\mathcal{I}$ was split into two disjoint subsets in a $1:1$ ratio, one used for training and the other reserved for evaluation. The “INJECTION" row of Table 5 demonstrates that CNL still achieves zero forgetting because the loss on Mastered Set can be explicitly calculated.
Subsequently, the same procedure is applied to the Mastered Set $\mathcal{M}$, as reported in the “MASTERED” row of Table 5. Under this setting, the loss on the evaluation Mastered subset cannot be directly computed and is instead estimated using the training Mastered subset. Due to this estimation error, CNL cannot guarantee a monotonic decrease in the evaluation loss. Nevertheless, while zero forgetting cannot be fully achieved, CNL still substantially alleviates catastrophic forgetting, reducing it by 59.1%–81.7% compared to conventional FT.
Furthermore, the training Mastered subset was combined with Injection Set during training to mitigate forgetting, a strategy known as replay. RP follows the same training setup as FT. Results in the “RP" row of Table 5 demonstrate that CNL not only learns more new knowledge but also forgets less mastered knowledge compared to replay.
These results demonstrate that CNL not only achieves zero forgetting under the ideal in-set condition, but also significantly alleviates catastrophic forgetting under more practical out-of-set evaluation settings.
@@ 6 Related Work
Table 5: Performance evaluation under out-of-set setting. Qwen2.5 1.5B is selected for experiment. To assess generalization, both Injection Set and Mastered Set were partitioned into training and evaluation subsets. This table reports the learning and forgetting performance on unseen evaluation data, demonstrating the excellent generalization capability of CNL.
Method MMLU MedQA ARC-C CSQA
Learned Forgot Learned Forgot Learned Forgot Learned Forgot
Injection FT 24 48 34 75 29 71 22 114
(7.3%) (5.5%) (9.7%) (13.2%) (17.9%) (7.5%) (16.2%) (13.8%)
CNL 31 0 39 0 29 0 23 0
(9.5%) (0.0%) (11.1%) (0.0%) (17.9%) (0.0%) (16.9%) (0.0%)
Mastered FT 178 110 208 143 104 49 86 94
(26.3%) (26.0%) (29.5%) (50.4%) (32.1%) (10.4%) (31.6%) (22.8%)
RP 157 61 124 58 86 14 70 28
(23.2%) (14.4%) (17.6%) (20.4%) (26.5%) (3.0%) (25.7%) (6.8%)
CNL 179 46 243 33 116 9 147 18
(26.4%) (10.9%) (34.5%) (11.6%) (35.8%) (1.9%) (54.0%) (4.4%)
@@ 6.1 Knowledge Injection
Existing knowledge injection methods are broadly categorized into two types: parametric methods and non-parametric methods Doan et al. (2021); Kotha et al. (2023).
Parametric methods directly modify model parameters to encode target knowledge via instruction tuning or continued pre-training, but often suffer from catastrophic forgetting of previously acquired knowledge Singhal et al. (2023); Han et al. (2023); Qiu et al. (2024). Non-parametric methods, such as retrieval-augmented generation, augment model inference with externally retrieved information without directly modifying model parameters Lewis et al. (2020); Ram et al. (2023); Xiong et al. (2024). As a result, they mitigate parameter-level interference, but their effectiveness is constrained by retrieval quality and limited access to implicit knowledge required for complex reasoning Salemi and Zamani (2024); Zhu et al. (2024); Kim et al. (2025) .
@@ 6.2 Mitigating Catastrophic Forgetting
Approaches to mitigate catastrophic forgetting have been extensively studied De Lange et al. (2021). Existing strategies generally operate along three dimensions.
Data-based approaches, such as experience replay, mix mastered knowledge with injected knowledge to preserve historical gradient signals, but suffer from storage overhead and privacy concerns. Constraint-based approaches, including PEFT, reduce forgetting by restricting the scope of parameter updates or penalizing large deviations from pretrained weights Wang et al. (2021); Hu et al. (2022). Localization-based approaches, such as knowledge editing methods, attempt to identify and modify task-relevant representation regions to achieve targeted knowledge updates with minimal side effects Meng et al. (2022, 20222).
Although effective in practice, most existing methods rely on heuristic designs and lack an explicit characterization of when and why forgetting occurs. In contrast, this work seeks to establish a principled connection between forgetting behavior and optimization dynamics.
6.3 Gradient Analysis in Continual Learning
From an optimization perspective, the gradient inner product (or gradient similarity) serves as a fundamental indicator of interference between learning objectives Yu et al. (2020); Liu et al. (2021). When gradients induced by new knowledge conflict with those associated with previously learned knowledge, optimization inevitably degrades prior performance, manifesting as catastrophic forgetting Farajtabar et al. (2020). Previous studies in continual and lifelong learning have leveraged gradient projection or alignment techniques to mitigate such conflicts at a global level Wang et al. (2025).
This work advances this perspective by decomposing gradient inner products at the neuron level, enabling identification of conflicting and collaborative components within the model. By selectively suppressing neuron-level conflicting gradients while preserving collaborative ones, forgetting mitigation is transformed from heuristic data balancing into a theoretically grounded training rule that is compatible with standard optimizers.
@@ Conclusion
Catastrophic forgetting during knowledge injection poses a fundamental challenge to the continual learning capability of LLMs. In this work, we establish a gradient-based theoretical framework to explain the underlying mechanism of catastrophic forgetting. We show that strongly negative gradient similarity is the fundamental cause of forgetting and further leverage this insight to categorize neurons into two types: conflicting neurons, which induce forgetting and account for 50%–75% of all neurons, and collaborative neurons, which mitigate forgetting and account for the remaining 25%–50%.
Based on this analysis, we propose a novel knowledge injection method, CNL, which freezes conflicting neurons and updates only collaborative neurons. We theoretically show that, under the assumption of an infinitesimal learning rate $\eta$ and exactly known Mastered Set, CNL prevents catastrophic forgetting. Extensive experiments on five LLMs, four datasets, and four optimizers further validate our analysis. CNL achieves zero forgetting under in-set evaluation and reduces catastrophic forgetting by 59.1%–81.7% under more practical out-of-set settings.
Overall, this work provides both theoretical insights and practical solutions for mitigating catastrophic forgetting in knowledge injection, offering a principled foundation for continual knowledge acquisition in LLMs.
Acknowledgments
This work was supported by the Noncommunicable Chronic Diseases-National Science and Technology Major Project under Grants 2024ZD0522702 and 2024ZD0528306, and the National Natural Science Foundation of China under Grant 62433007.
@@ Appendix A Dataset Details
To verify our conclusions, we selected four datasets for experimentation: MMLU, MedQA, ARC-C, and CSQA. The data processing method is described as follows:
  • MMLU: 1/10 subsampling is performed to result in 1, 555 remaining questions.
  • MedQA: 1/10 subsampling is performed to result in 1, 273 remaining questions.
  • ARC-C: 1/2 subsampling is performed to result in 1, 270 remaining questions.
  • CSQA: 1/10 subsampling is performed to result in 1, 096 remaining questions.

### v2

@@ 5.1 Establishment of CPL
Since Conflicting Parameters are responsible for forgetting, a natural question arises: can forgetting be mitigated by freezing Conflicting Parameters and training only Collaborative Parameters? We refer to this learning paradigm as Collaborative Parameter Learning (CPL).
To indicate the sign of gradient similarity for the $j$-th parameter, we define indicator function $I(\cdot)$ as
@@
Applying the indicator function to all parameter yields
$\mathbb{I}[S(\mathcal{M},\mathcal{I})\geq 0]=\big(I(s_{\theta_{1}}(\mathcal{M},\mathcal{I})\geq 0),\ldots,I(s_{\theta_{|\theta|}}(\mathcal{M},\mathcal{I})\geq 0)\big)^{\top}$ (15)
CPL freezes parameter with negative gradient similarity, leading to a parameter update rule defined as
@@
where $\odot$ denotes the Hadamard product. The change in loss on $\mathcal{M}$ after parameter update is given by
$\Delta L^{\mathcal{M}}(\theta)$ $=-\eta\,\nabla_{\theta}L^{\mathcal{M}}(\theta)^{\top}\big(\mathbb{I}[S(\mathcal{M},\mathcal{I})\geq 0]\odot\nabla_{\theta}L^{\mathcal{I}}(\theta)\big)$
@@
This indicates that, given full access to the samples in $\mathcal{M}$, under a first-order approximation and with an infinitesimally small learning rate, CPL guarantees that the loss on $\mathcal{M}$ does not increase during knowledge injection.
To experimentally evaluate the effectiveness of CPL, we performed knowledge injection using both CPL and conventional fine tuning (FT) under an ideal situation where $\mathcal{M}$ can be fully accessed. The case where $\mathcal{M}$ is not explicitly available will be discussed in Section 6.3.
CPL and FT are compared under the premise that CPL learns no fewer questions than FT. Table 3 shows that FT forgets a number of mastered samples comparable to the amount of newly injected knowledge, whereas CPL forgets zero samples when $\mathcal{M}$ is fully available. This demonstrates that CPL effectively mitigates forgetting while maintaining strong capability for knowledge injection.
5.2 Comparison with Baseline Methods
Table 4: Comparison with baseline methods on Qwen2.5 1.5B. Results are reported in the format of “Number of questions (Percentage)”. CPL maintains stronger capability both for knowledge injection and forgetting mitigating.
MMLU MedQA ARC-C CSQA
@@ 5.2 Comparison with Baseline Methods
EWC [10] 164 178 165 209 87 163 97 70
(24.2%) (20.3%) (23.4%) (36.8%) (26.9%) (17.2%) (35.7%) (8.5%)
MAS [11] 167 200 184 259 104 80 86 161
(24.7%) (22.8%) (26.1%) (45.6%) (32.1%) (8.5%) (31.6%) (19.6%)
MOFO [29] 167 178 193 268 103 92 87 165
(24.7%) (20.3%) (27.4%) (47.2%) (31.8%) (9.7%) (32.0%) (20.0%)
WSC [30] 115 97 113 116 76 44 68 96
(17.0%) (11.0%) (16.1%) (20.4%) (23.5%) (4.7%) (25.0%) (11.7%)
GPM [18] 165 153 190 257 102 73 83 120
(24.4%) (17.4%) (27.0%) (45.2%) (31.5%) (7.7%) (30.5%) (14.6%)
OGD [16] 128 6 142 1 110 3 84 1
(18.9%) (0.7%) (20.2%) (0.2%) (34.0%) (0.3%) (30.9%) (0.1%)
A-GEM [17] 129 4 155 1 110 3 84 1
(19.1%) (0.5%) (22.0%) (0.2%) (34.0%) (0.3%) (30.9%) (0.1%)
CPL 168 0 193 0 163 0 147 0
(24.8%) (0.0%) (27.4%) (0.0%) (50.3%) (0.0%) (54.0%) (0.0%)
Table 5: Comparison of computational cost on Qwen2.5 1.5B. CPL replaces the high-precision vector projection with low-precision element-wise sign comparison, reducing peak GPU VRAM by approximately 4.3 GB and time per epoch by 16.5% compared to A-GEM.
Peak GPU VRAM (G) Time / Epoch (s)
EWC 29.5 3867
MAS 31.9 4836
MOFO 20.1 2170
WSC 53.0 1808
GPM 14.9 7504
OGD 20.6 3738
A-GEM 20.6 3508
CPL 16.3 ($\downarrow$ 4.3 GB) 2928 ($\downarrow$ 16.5%)
We further compare CPL with seven baseline methods: EWC [10], MAS [11], MOFO [29] and WSC [30], GPM [18], OGD [16], and A-GEM [17]. Table 4 demonstrates that gradient-based methods (OGD and A-GEM) exhibit significantly stronger forgetting mitigation than the other baseline methods. Moreover, they are also more conceptually similar to CPL. Therefore, we focus on the comparison between CPL and gradient-based approaches.
In terms of forgetting mitigation, CPL slightly outperforms gradient-based methods, and significantly surpasses other methods. In terms of knowledge acquisition, unlike traditional gradient-based methods, which require the gradients of $\mathcal{M}$ and $\mathcal{I}$ to be completely orthogonal, CPL masks only the conflicting component while preserving collaborative component, resulting 20.2%-48.2% more injected samples than OGD and A-GEM within the specified 25 epochs.
The computational cost is compared in Table 5. Gradient-methods require high-cost vector storage for orthogonalization (FP32, accounts for 4 GB per billion parameters), whereas CPL performs element-wise sign comparison and only requires low-costs vector storage (INT8, accounts for 1 GB per billion parameters). Therefore, CPL reduces peak GPU VRAM by approximately 3 GB per billion parameters, which is confirmed by the 4.3 GB GPU VRAM reduction observed in Table 5 for Qwen2.5 1.5B. Additionally, the simplification from high-cost vector projection to element-wise sign comparison leads to a 16.5% reduction in computational overhead (from 3508 s per epoch for OGD to 2928 s per epoch for CPL).
5.3 Collaborative Parameters Consumption Analysis
Since CPL freezes Conflicting Parameters and only updates Collaborative Parameters, it is important to learn that whether Collaborative Parameters are adequately preserved across different datasets. To answer this question, the change in Collaborative Parameters for each dataset is recorded after training on a specific dataset.
Table 6 shows that Collaborative Parameters are task-dependent. After training on a dataset, the Collaborative Parameters for that dataset decrease, but the Collaborative Parameters for the other datasets even increase. This indicates that CPL does not simply consume a fixed global pool of parameters; instead, the set of Collaborative Parameters changes with the injected knowledge.
Moreover, CPL is conducted on larger-scale datasets to investigate whether Collaborative Parameters are sufficient for knowledge injection. In Appendix G.1, joint multi-dataset learning is conducted. Four datasets are mixed together and trained simultaneously. Table 10 shows that FT causes 91.7% forgetting on mixed dataset, while CPL forgets 0.0% questions. In Appendix G.2, four datasets are trained sequentially to conduct sequential multi-dataset learning. Table 11 shows that CPL nearly do not forget mastered questions across all learning orders, while FT suffers significant forgetting.
Additionally, the learning dynamics of CPL and FT are analyzed in Appendix H. Although CPL freezes Conflicting Parameters and only updates Collaborative Parameters, the speed of knowledge injection for CPL is not noticeably slowed than conventionally FT, demonstrating adequate Collaborative Parameters for training.
6 Extending CPL to Diverse Scenarios
6.1 Extending CPL to More Optimizers
In Appendix I, we extend CPL to a broader family of optimizers, including Momentum, Adam, and AdamW. Experimental results in Table 12 prove CPL forgets zero questions across different optimizers under the ideal situation that $\mathcal{M}$ is available. Moreover, CPL is also theoretically extended to arbitrary optimizers in Appendix I.4.
6.2 Extending CPL to Various Tasks
In Appendix J, CPL is further extended to diverse tasks including: sentiment classification (Yahoo Answers, AG News, and DBPedia) [23], open-ended question answering (NumGLUE-cm and NumGLUE-ds) [24], multilingual settings (C-STANCE) [25], and multimodal tasks (MMBench) [26]. Table 13 demonstrates that CPL effectively mitigates forgetting across diverse tasks.
6.3 Extending CPL to Out-of-Set Setting
Table 6: Change in Collaborative Parameters after training. The dataset in row is the training dataset, and the dataset in column is the evaluation dataset. After training on a specific dataset, its own Collaborative Parameters decrease, but Collaborative Parameters for the others do not decrease.
DATASET MMLU MedQA ARC-C CSQA
MMLU -12.1% $20.8\% +$$14.3\% +$$13.3\%$
MedQA $8.8\% -$8.2% $1.1\% +11$$1.0\%$
ARC-C $13.9\% +$$17.6\% -$10.5% $14.1\%$
CSQA $12.4\% +$$23.7\% +$$10.4\% -$9.6%
In the previous sections, we conducted experiments under an ideal setting that all samples in $\mathcal{M}$ are accessible. Herein, CPL is extended to out-of-set generalization settings where only a subset of $\mathcal{M}$ is accessible in Appendix K. As shown in Table 14, when $\mathcal{M}$ is partially accessible, although CPL forgets more question than the ideal setting, it still mitigates forgetting by 59.1%–81.7% than conventional FT. The result also indicates that incorporating the available subset using CPL is more effective in forgetting mitigation than using replay.
6.4 Extending CPL to Cross-Prompt Setting
To study CPL mereley injects specific prompts or the underlying knowledge, four prompts are designed for each question as shown in Appendix L. CPL is trained on two prompts and evaluated on the other two prompts. Table 15 shows that CPL significantly mitigates forgetting on unseen prompts.
@@ 7.1 Knowledge Injection
Knowledge injection methods are commonly divided into parametric and non-parametric approaches [31, 32]. Parametric methods, such as instruction tuning, update model weights, but can overwrite previously acquired knowledge [33, 34, 35]. Non-parametric methods, such as retrieval-augmented generation, avoid parameter updates by using external evidence at inference time [3, 36, 37], but remain limited by retrieval quality [38, 39, 40].
@@ 7.2 Mitigating Catastrophic Forgetting
For model adaptation, forgetting is typically mitigated through replay, constraints, localization, or update selection [9]. Replay mixes old and new data but introduces storage and privacy costs [41]; constraint methods such as PEFT, EWC, and MAS restrict updates or protect important weights [12, 13, 10, 11]; localization and editing methods target task-relevant regions [42, 43]; and recent methods such as MOFO and WSC select or consolidate updates during adaptation [29, 30]. Gradient similarity provides a direct signal of interference between learning objectives [14, 15]: conflicting gradients can degrade prior knowledge and induce forgetting [16]. Prior projection or alignment methods, including OGD, A-GEM, and GPM, mitigate such conflicts at a global level [16, 17, 18, 44].
Limitations
First, our analysis relies on a first-order approximation. However, LLMs are generally non-convex. Future work could explore non-linear dynamics of forgetting. Second, CPL assumes access to $\mathcal{M}$. Although we show that CPL can still mitigate forgetting when only a subset of $\mathcal{M}$ is accessible, it remains an open question how to best estimate gradient similarity without direct access to $\mathcal{M}$. Future work could explore proxy methods for gradient estimation. Third, the gradient computation of $\mathcal{M}$ requires additional computational cost. Although CPL reduces computational cost compared with other existing gradient-based methods, it still incurs more cost than conventional FT. Future work could explore more efficient approximations of gradient similarity.
@@ Conclusion
Catastrophic forgetting during knowledge injection prevents LLMs from retaining previously learned knowledge. We show that strongly negative gradient similarity drives forgetting, and categorize parameters into Conflicting and Collaborative Parameters. CPL freezes Conflicting Parameters and updates only Collaborative Parameters. Experiments show that CPL significantly mitigates forgetting while maintaining knowledge injection capability across diverse scenarios.
Impact Statement
This study provides parameter-level understanding for forgetting during knowledge injection and reveals the role of Conflicting and Collaborative parameters. Building on this insight, CPL is proposed by freezing Conflicting Parameters and updating only Collaborative Parameters. Like other knowledge injection methods, CPL may inject incorrect, biased, or harmful knowledge if the injected knowledge is poorly curated, so practical deployment requires careful data governance.
@@ Appendix A Dataset Details
MMLU, MedQA, ARC-C, and CSQA are employed as foundation datasets, which are 1/10, 1/10, 1/2, and 1/10 randomly subsampled from the original datasets, respectively, resulting in 1,555, 1,273, 1,270, and 1,096 questions, respectively.
Yahoo Answers, AG News, DBPedia, NumGLUE-cm, NumGLUE-ds, C-STANCE, and MMBench are used for extensive study. All datasets are randomly subsampled to 500 questions each, except for MMBench, which contains 500 Chinese and 500 English questions.